# Contrastive Learning Augmented Social Recommendations

### v2

@@ Abstract.
Recommender systems play a pivotal role in modern content platforms, yet traditional behavior-based models often face challenges in addressing cold users with sparse interaction data. Engaging these users, however, remains critical for sustaining platform growth. To tackle this issue, we propose leveraging reconstructed social graph to complement interest representations derived from behavioral data. Despite the widespread availability of social graphs on content platforms, their utility is hindered by social-relation noise and inconsistencies between social and behavioral interests. To mitigate noise propagation in graph data and extract reliable social interests, we introduce a dual-view denoising framework. This approach first applies low-rank singular value decomposition (SVD) to the user-item interaction matrix, generating denoised user embeddings for reconstructing the social graph. It then employs contrastive learning to align the original and reconstructed social graphs. To address the discrepancy between social and behavioral interests, we utilize a mutual distillation mechanism that decomposes interests into four subcategories: aligned social/behavioral interests and social/behavioral-specific interests, enabling effective integration of the two. Empirical results demonstrate the efficacy of our method, particularly in improving recommendations for cold users, by combining social and behavioral data. The implementation of our approach is publicly available at https://github.com/WANGLin0126/CLSRec.
@@
*Corresponding author.
submissionid: 1638copyright: noneccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems (RS) have now become an indispensable component of modern content platforms (e.g., music, video, review&rating), which provide personalized content to shape user experiences and drive engagement. Traditional recommendation models(cheng2016wide, ; zhou2018deep, ; li2023stan, ; ouyang2023click, ; lin2019pareto, ), and typical content recommendation models(covington2016deep, ; xiao2024lottery4cvr, ; xiao2023neighbor, ; zhang2023modeling, ; ouyang2023knowledgesoftintegrationmultimodal, ), typically score candidate items based on user characteristics extracted from behaviors (e.g., clicking, viewing). However, behavior-based recommendation models typically fail on cold users, who exhibit little behavior or are new to the platform (gope2017survey, ; sethi2021cold, ; bobadilla2012collaborative, ; lu2020meta, ; zhang2019contextual, ; li2017user, ). Improving the personalized experience of cold users on RS can be helpful to drive engagement, and converting cold users into active users is vital to the continuous development of all content platforms.
To overcome the limitations of behavior-based models, we propose to leverage the social graph to complete the interest patterns. Social graphs on popular content platforms, including LastFM, TikTok, and YouTube, are prevalent and easily accessible. Typical social graphs encompasse a variety of user relationships, such as mutual friendships, shared interests in movie stars, and subscriptions to the same TV series (Fig.1). It stands to reason that users who engage in these social interactions may share similar interest preferences. Consequently, integrating the social graph into RS could yield significant benefits. Nevertheless, despite the wealth of information in social graphs, behavior-based models face challenges in affectively harnessing this data.
@@
Social Relation Noise. The heterogeneity of graph data introduces noise propagation, which makes the social interest inaccurate (xiao2023incorporating, ). For example, users who are friends tend to have higher similarity, and their information propagation is considered beneficial for he performance of the recommendation system. However, the information propagation between celebrities and their fans needs to be carefully considered to avoid degrading the performance. On the other hand, considering the sparsity of social data, there is a significant proportion of users who have limited social connections. To enhance the representation capability of this sub-data in social relation, we adopt a dual-view denoising approach. The approach firstly applies low-rank SVD to the user-item matrix to obtain user embeddings. The advantage of low-rank decomposition is the ability to select principal components and remove unnecessary interference, effectively filtering the user-item matrix. Then, the user embeddings are used to calculate the similarity between all pairs of users for a reconstructed social graph. The contrastive learning between these two graphs is finally adopted as data augmentation for users, alleviating the challenges imposed by graph noise.
@@
Interest Inconsistency. Although social relations provide numerous information for each user, there exists distinct interest inconsistency between the social interest representation and the behavioral interest representation (xiao2023incorporating, ), which exposes challenge to effectively fuse the knowledge existing in the two interest. We adopt a ”mutual distillation” method to co-extract the interest representations of social graphs and interactive behaviors. From the view of social interest, we distill the information from the social graph that is useful to the user interest representation. On the other hand, from the view of behavior interest, to avoid the knowledge loss in the behaviors, we isolate two distinct interest representations from the behavior-based interest, i.e., social-interest similar representation and behavior-specific representation. In this way, we fully fuse the knowledge of two domains for recommendation.
@@
  • We propose a novel social graph-enhanced method to traditional behavior-based graph recommendations, which incorporates social-relation de-noising and social-behavior alignment, and comprehensively exploit the knowledge in social relations.
  • Our method provides a dual-view contrast learning to purify the raw social relations and uses ”mutual distillation” to co-extract the information from behaviors and social relations, which maximally fuse knowledge in both two interest representations without significant knowledge loss.
  • Experimental results on widely adopted industry datasets demonstrate the effectiveness of our method, especially for cold users, verifying that effectively fusing social relations and behaviors can be highly useful for modern recommendation platforms.
@@ 2. Methodology
In this section, we present our proposed Contrastive Learning Social Recommendation (CLSRec). It employs GNN to represent user-item behavior interaction and then filters the interaction graph to remove invalid connections, reconstructing a refined social graph view as the basis for generating refined user embeddings. To integrate user embeddings from the reconstructed and original graphs, we employ contrastive learning, which captures complementary information from both graphs. Given the two user representations derived from behaviors and social relations, we address the inconsistency between social interests and private behaviors. To achieve this, we disentangle the consistent and inconsistent information embedded in these representations. This disentanglement allows us to maximize the utility of the knowledge contained within both representations and construct a comprehensive user embedding.
@@
2.1. Scoial Recommendation
Social recommendation approaches leverage two primary types of graphs to derive user and item embeddings: the user-item interaction graph and the user-user social graph. Specifically, the propagation for user-item interaction graph is modeled as follows:
@@ 2.2. Refine the Social knowledge
Social graphs are valuable in social recommendation as they can reflect purchasing behaviors, such as friends with strong ties sharing similar product preferences. However, they often include invalid relationships, such as users with limited social interactions, which lack mutual interests and can degrade recommendation performance. To address this, we propose reconstructing a refined social graph view using embeddings derived from the interaction graph through Singular Value Decomposition (SVD). Subsequently, contrastive learning is employed to align the user embeddings obtained from the reconstructed graph with those derived from the original social graph. This alignment process aims to suppress noise and enhance the performance of recommendations.
@@ 2.2.1. Reconstructed Social View
To construct the refined social graph view, SVD (abdi2007singular, ; cai2023lightgcl, ) is first utilized to decompose the user-item interaction matrix $A\in\mathbb{R}^{M\times N}$ as $A=USV^{\top}$. Here, $S\in\mathbb{R}^{M\times N}$ is a diagonal matrix of singular values, while $U\in\mathbb{R}^{M\times M}$ and $V\in\mathbb{R}^{N\times N}$ are orthogonal matrices of eigenvectors. The dimensions $M$ and $N$ represent the number of users and items, respectively. Singular values reflect the significance of their corresponding eigenvectors, referred to as principal components. To construct a low-rank approximation, we select the top $k$ singular values and their associated components, forming:
(4) $A_{k}=U_{k}S_{k}V_{k}^{\top},$
where $A_{k}$ is a rank-$k$ approximation, $U_{k}\in\mathbb{R}^{M\times k}$ and $V_{k}\in\mathbb{R}^{N\times k}$ contain the top $k$ eigenvectors, and $S_{k}\in\mathbb{R}^{k\times k}$ is a diagonal matrix of the $k$ largest singular values.
Then, a refined social view is reconstructed as,
@@ 2.3. Interest Inconsistency Alignment
Although social relations and user behaviors both contribute to user interest representation, inconsistencies often arise between a user’s social interests and private behaviors. Directly combining these representations from different data domains may introduce noise and reduce effectiveness. To address this, we propose a ”mutual distillation” method to extract valuable information from social relations and behaviors while isolating conflicting information to mitigate noise (Fig. 3).
For interest alignment, we utilize a co-attention mechanism to differentiate between shared and distinct interests in the representations of behaviors ($E_{u}$) and social relations ($\hat{E}_{u}$). Through the mutual distillation process, we derive bivariate weight vectors that quantify the alignment of each dimension in the two interest vectors. Using these weights, each representation is decomposed into two components: aligned interests and specific interests.
@@
where $W_{b},W_{s}\in R^{d\times d}$ are learnable weight matrices, and $softmax$ normalizes the weights. The aligned weights $W_{b}$ and $W_{s}$ represent affinity scores for behavior and social interests, respectively. Elements in $E_{u}$ and $\hat{E}_{u}$ with higher weights in $W_{b}$ and $W_{s}$ are indicative of aligned interests, while those with lower weights capture specific interests. The isolating process is defined as:
(10) $E_{ua}$ =$=\{E_{uj}\mid W_{bj}\geq\gamma_{i}\},\quad E_{uu}=\{E_{uj}\mid W_{bj}<\gamma_{i}\},$
(11) $\hat{E}_{ua}$ =$=\{\hat{E}_{uj}\mid W_{sj}\geq\gamma_{s}\},\quad\hat{E}_{uu}=\{\hat{E}_{uj}\mid W_{sj}<\gamma_{s}\},$
@@
This process enables users to retain both popular interests shared among social relations and individual-specific interests. As a result, we derive three representations: $E_{u},E_{ua},E_{uu}$ from the behavioral perspective, and $\hat{E}_{u},\hat{E}_{ua},\hat{E}_{uu}$ from the social perspective. To integrate these representations, we employ a gated routing strategy. The gate input is a statistical feature vector, $user\_fea$, capturing the frequency of behavior and the diversity of social relations. This allows users to dynamically emphasize the most expressive representation based on the richness of their interest types. The dynamic routing operation is defined as:
@@
Table 1. Recommendation performance
@@
BPR DiffNet LightGCN SocialLGN
@@
Improve BPR DiffNet LightGCN SocialLGN
@@
LastFM Precision@10 0.0922 0.1727 0.1961 0.1972 0.1997 1.29% 0.0282 0.0417 0.0417 0.0458 0.0542 18.27%
Precision@20 0.0720 0.1215 0.1358 0.1368 0.1381 0.92% 0.0209 0.0271 0.0313 0.0333 0.0396 18.87%
Recall@10 0.0962 0.1779 0.2003 0.2026 0.2045 0.94% 0.1151 0.1713 0.1727 0.1974 0.2245 13.75%
Recall@20 0.1499 0.2488 0.2769 0.2794 0.2821 0.95% 0.1615 0.2407 0.2416 0.2663 0.2917 9.55%
NDCG@10 0.1099 0.2219 0.2536 0.2566 0.2615 1.92% 0.0828 0.1107 0.1374 0.1419 0.1573 10.88%
NDCG@20 0.1321 0.2474 0.2788 0.2883 0.2869 -0.47% 0.0989 0.1309 0.1560 0.1643 0.1813 10.33%
Ciao Precision@10 0.0145 0.0238 0.0271 0.0276 0.0284 2.81% 0.0061 0.0104 0.0131 0.0134 0.0134 0.03%
Precision@20 0.0111 0.0182 0.0202 0.0205 0.0208 1.36% 0.0047 0.0081 0.0096 0.0097 0.0099 2.14%
Recall@10 0.0220 0.0341 0.0410 0.0430 0.0425 -1.10% 0.0208 0.0339 0.0429 0.0441 0.0436 -1.14%
Recall@20 0.0339 0.0527 0.0591 0.0618 0.0623 0.78% 0.0328 0.0539 0.0616 0.0630 0.0636 1.03%
NDCG@10 0.0229 0.0359 0.0437 0.0441 0.0452 2.43% 0.0138 0.0248 0.0319 0.0328 0.0331 0.98%
NDCG@20 0.0260 0.0403 0.0478 0.0486 0.0494 1.66% 0.0179 0.0316 0.0384 0.0394 0.0401 1.74%
@@ 2.4. Training and Optimization
(14) $\mathcal{L}_{bpr}=-\sum_{u=1}^{M}\sum_{i\in\mathcal{N}_{u}}\sum_{j\notin\mathcal{N}_{u}}\mathrm{ln}\sigma(\hat{y}_{ui}-\hat{y}_{uj})+\lambda||E^{(0)}||^{2},$
@@ 3. Experiments
Our experiments are conducted on two real-world industry datasets, which vary in terms of scale and sparsity: LastFM [^1] https://grouplens.org/datasets/hetrec-2011/ and Ciao [^2] https://www.cse.msu.edu/ tangjili/datasetcode/truststudy.htm. LastFM is a music-related dataset that records the friendships of 1,892 users and their interactions with 17,632 music items. Ciao is a database covering online shopping, containing ratings from 7,375 users on 105,114 products as well as the social relationships between users. During the experiments, we randomly split the interaction data of each dataset into a training set and a test set, with 80% of the data used for training and the remaining 20% used for testing. To fairly tune the hyperparameters of the model and prevent overfitting, we randomly select 10% of the training data as a validation set.
To evaluate the performance of CLSRec, we compare it with the following 4 state-of-the-art methods: BPR(rendle2012bpr, ), a classic pairwise collaborative filtering algorithm optimized by maximizing the difference between the probability of a user interacting with positive samples and negative samples. DiffNet(wu2019neural, ), a graph-based social recommendation method that generates final user representations by integrating user representation vectors from both the user-item interaction graph and the social graph. LightGCN(he2020lightgcn, ), a widely applied model based on NGCF that further simplifies the model structure. SocialLGN(liao2022sociallgn, ), that is currently the most advanced social recommendation graph model based on LightGCN.

### v3

# Social Relation Meets Recommendation: Augmentation and Alignment

@@ Abstract.
Recommender systems are essential for modern content platforms, yet traditional behavior-based models often struggle with cold users who have limited interaction data. Engaging these users is crucial for platform growth. To bridge this gap, we propose leveraging the social-relation graph to enrich interest representations from behavior-based models. However, extracting value from social graphs is challenging due to relation noise and cross-domain inconsistency. To address the noise propagation and obtain accurate social interest, we employ a dual-view denoising strategy, employing low-rank SVD to the user-item interaction matrix for a denoised social graph and contrastive learning to align the original and reconstructed social graphs. Addressing the interest inconsistency between social and behavioral interests, we adopt a ”mutual distillation” technique to isolate the original interests into aligned social/behavior interests and social/behavior specific interests, maximizing the utility of both. Experimental results on widely adopted industry datasets verify the method’s effectiveness, particularly for cold users, offering a fresh perspective for future research. The implementation can be accessed at https://github.com/WANGLin0126/CLSRec.
@@
*Corresponding author
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.376087ccs: Information systems Recommender systemsccs: Information systems Social recommendation
@@ 1. Introduction
Recommender systems (RS) have now become core parts of major Internet platforms, providing personalized content to improve the user experience and drive engagement. Traditional recommendation models(cheng2016wide, ; zhou2018deep, ; li2023stan, ; ouyang2023click, ; lin2019pareto, ), and typical content recommendation models(covington2016deep, ; xiao2024lottery4cvr, ; xiao2023neighbor, ; zhang2023modeling, ; ouyang2023knowledgesoftintegrationmultimodal, ), score candidate items based on user characteristics extracted from behaviors (e.g., clicking, viewing). However, behavior-based recommendation models typically fail on cold users, who exhibit little behavior or are new to the platform (gope2017survey, ; sethi2021cold, ; bobadilla2012collaborative, ; lu2020meta, ; zhang2019contextual, ; li2017user, ). Improving the personalized experience of cold users on RS can be helpful to drive engagement, and converting cold users into active users is vital to the continuous development of all content platforms.
We propose leveraging social graphs for joint social-behavior modeling, shifting beyond behavior-centric approaches. Social graphs capture diverse user relations—friendships, shared interests, and subscriptions—across platforms like YouTube, LastFM, and TikTok. Such relations often imply similar preferences, making social graph integration highly beneficial for recommendation systems. However, behavior-based models still struggle to effectively utilize this rich information.
@@
Social Relation Noise. The heterogeneity of graph data introduces noise propagation, making the social interest inaccurate (xiao2023incorporating, ). For example, users who are friends tend to have higher similarity, and their information propagation is considered beneficial for the performance of recommendation systems. However, the information propagation between celebrities and their fans needs to be carefully considered to avoid degrading the performance. On the other hand, considering the sparsity of social data, there is a significant proportion of users who have limited social connections. To enhance the representation capability of this sub-data in social relation, we adopt a dual-view denoising approach. The approach firstpplies low-rank SVD to the user-item matrix to obtain denoised user embeddings for a reconstructed social graph. Then, contrastive learning between original and reconstructed social graphs is adopted as data augmentation for users, alleviating graph noise.
@@
Interest Inconsistency. Although social relations provide numerous information for each user, there exists distinct interest inconsistency between the social interest representation and the behavioral interest representation (xiao2023incorporating, ), which exposes challenge to effectively fuse the knowledge existing in the two interest. We adopt a ”mutual distillation” method to co-extract the interest representations of social graphs and interactive behaviors. From the view of social interest, we distill the information from the social graph that is useful to the user-interest representation. From the view of behavior interest, we isolate two distinct interest representations from the behavior-based interest, i.e., social-interest similar representation and behavior-specific representation, to avoid knowledge loss in behaviors. In this way, we fully fuse the knowledge of two domains for recommendation.
@@
  • We introduce an innovative social graph augmentation method that enhances conventional behavior-driven recommendation models through holistic user interest modeling.
  • The method employs dual-view contrast learning to denoise original social relations and ”mutual distillation” to co-extract knowledge from both behaviors and social relations for maximal knowledge fusing, offering a promising fresh paradigm.
  • Experiments on widely adopted industry datasets demonstrate the effectiveness of the approach, particularly for cold users, highlighting its potential for broader adoption in social-behavior recommendation.
@@ 2. Methodology
In this section, we present our proposed Contrastive Learning Social Recommendation (CLSRec). CLSRec first employs GNNs to learn user–item interaction embeddings, then filters noisy connections to reconstruct a denoised social graph for refined user representations. To enrich these representations, contrastive learning integrates embeddings from both the original and reconstructed graphs, capturing complementary information. Finally, CLSRec disentangles consistent and inconsistent signals between social interests and private behaviors, maximizing the utility of both for comprehensive user modeling.
@@
2.1. Information Extraction of Dual-view
Social recommendation leverage two primary types of graphs to derive user and item embeddings: the user-item interaction graph and the user-user social graph. Specifically, the propagation for user-item interaction graph is modeled as follows:
@@ 2.2. Refine the Social knowledge
Social graphs capture interest similarity (e.g., friends with shared preferences) but also contain noisy or weak ties that harm recommendation quality. To mitigate this, we reconstruct a refined social graph using user embeddings derived from the interaction graph via Singular Value Decomposition (SVD). Contrastive learning is then applied to align embeddings from the original and reconstructed graphs, improving recommendation performance.
@@ 2.2.1. Reconstructed Social View
To construct the refined social graph view, SVD (abdi2007singular, ; cai2023lightgcl, ) is first utilized to decompose the user-item interaction matrix $A\in\mathbb{R}^{M\times N}$ as $A=USV^{\top}$. Here, $S\in\mathbb{R}^{M\times N}$ is a diagonal matrix of singular values, while $U\in\mathbb{R}^{M\times M}$ and $V\in\mathbb{R}^{N\times N}$ are orthogonal matrices of eigenvectors. The dimensions $M$ and $N$ represent the number of users and items, respectively. Singular values reflect the significance of their corresponding eigenvectors, referred to as principal components. To construct a low-rank approximation, we select the top $k$ singular values and their associated components, forming: $A_{k}=U_{k}S_{k}V_{k}^{\top}$, where $A_{k}$ is a rank-$k$ approximation, $U_{k}\in\mathbb{R}^{M\times k}$ and $V_{k}\in\mathbb{R}^{N\times k}$ contain the top $k$ eigenvectors, and $S_{k}\in\mathbb{R}^{k\times k}$ is a diagonal matrix of the $k$ largest singular values. Then, a refined social view is reconstructed as,
@@ 2.3. Interest Inconsistency Alignment
Although both social relations and user behaviors inform user interests, they may be inconsistent, and naive fusion risks introducing noise. To mitigate this, we propose a mutual distillation method that extracts complementary signals while isolating conflicts (Fig. 3). Specifically, a co-attention mechanism compares behavior embeddings ($E_{u}$) and social embeddings ($\hat{E}_{u}$), generating weight vectors to decompose each into aligned and specific interests.
@@
where $W_{b},W_{s}\in R^{d\times d}$ are learnable weight matrices, and softmaxsoftmax normalizes the weights. The aligned weights $W_{b}$ and $W_{s}$ represent affinity scores for behavior and social interests, respectively. Elements in $E_{u}$ and $\hat{E}_{u}$ with higher weights in $W_{b}$ and $W_{s}$ are indicative of aligned interests, while those with lower weights capture specific interests. The isolating process is defined as:
(9) $E_{ua}$ $=\{E_{uj}\mid W_{bj}\geq\gamma_{i}\},\quad E_{uu}=\{E_{uj}\mid W_{bj}<\gamma_{i}\},$
(10) $\hat{E}_{ua}$ $=\{\hat{E}_{uj}\mid W_{sj}\geq\gamma_{s}\},\quad\hat{E}_{uu}=\{\hat{E}_{uj}\mid W_{sj}<\gamma_{s}\},$
@@
This process preserves both shared and individual-specific interests, yielding three representations from behaviors ($E_{u},E_{ua},E_{uu}$) and socials ($\hat{E}u,\hat{E}{ua},\hat{E}_{uu}$). To fuse them, we adopt a gated routing strategy, where the gate input ($user_{f}ea$) encodes behavior frequency and social diversity, enabling users to dynamically emphasize the most informative representation. The routing is defined as:
@@
Table 1. Recommendation performance comparison. The best results are highlighted in bold. Improve denotes the relative improvement (%) of our method over the best baseline.
@@
BPR DiffNet LightGCN SocialLGN GBSR
@@
Improve BPR DiffNet LightGCN SocialLGN GBSR
@@
LastFM Precision@10 0.0922 0.1727 0.1961 0.1972 0.1979 0.2002 1.16% 0.0282 0.0417 0.0417 0.0458 0.0511 0.0583 14.09%
Precision@20 0.0720 0.1215 0.1358 0.1368 0.1370 0.1384 1.02% 0.0209 0.0271 0.0313 0.0333 0.0361 0.0417 15.51%
Recall@10 0.0962 0.1779 0.2003 0.2026 0.2030 0.2057 1.33% 0.1151 0.1713 0.1727 0.1974 0.2089 0.2353 12.64%
Recall@20 0.1499 0.2488 0.2769 0.2794 0.2800 0.2829 1.04% 0.1615 0.2407 0.2416 0.2663 0.2690 0.3056 13.61%
NDCG@10 0.1099 0.2219 0.2536 0.2566 0.2573 0.2617 1.71% 0.0828 0.1107 0.1374 0.1419 0.1510 0.1624 7.55%
NDCG@20 0.1321 0.2474 0.2788 0.2883 0.2885 0.2872 -0.45% 0.0989 0.1309 0.1560 0.1643 0.1730 0.1852 7.05%
Ciao Precision@10 0.0145 0.0238 0.0271 0.0276 0.0280 0.0285 1.79% 0.0061 0.0104 0.0131 0.0134 0.0135 0.0140 3.70%
Precision@20 0.0111 0.0182 0.0202 0.0205 0.0201 0.0212 3.41% 0.0047 0.0081 0.0096 0.0097 0.0096 0.0101 4.12%
Recall@10 0.0220 0.0341 0.0410 0.0430 0.0431 0.0441 2.32% 0.0208 0.0339 0.0429 0.0441 0.0445 0.0458 2.92%
Recall@20 0.0339 0.0527 0.0591 0.0618 0.0622 0.0635 2.09% 0.0328 0.0539 0.0616 0.0630 0.0638 0.0648 1.57%
NDCG@10 0.0229 0.0359 0.0437 0.0441 0.0448 0.0456 1.79% 0.0138 0.0248 0.0319 0.0328 0.0330 0.0337 2.12%
NDCG@20 0.0260 0.0403 0.0478 0.0486 0.0488 0.0502 2.87% 0.0179 0.0316 0.0384 0.0394 0.0401 0.0405 1.00%
@@ 2.4. Training and Optimization
[…]
@@ 3. Experiments
Our experiments are conducted on two real-world industry datasets, which vary in terms of scale and sparsity: LastFM [^1] https://grouplens.org/datasets/hetrec-2011/ and Ciao [^2] https://www.cse.msu.edu/ tangjili/datasetcode/truststudy.htm. LastFM is a music-related dataset that records the friendships of 1,892 users and their interactions with 17,632 music items. Ciao is a database covering online shopping, containing ratings from 7,375 users on 105,114 products as well as the social relationships between users. During the experiments, we randomly split the interaction data of each dataset into a training set and a test set, with 80% of the data used for training and the remaining 20% used for testing. To fairly tune the hyperparameters and prevent overfitting, we randomly select 10% of the training data as a validation set.
To evaluate the performance of CLSRec, we compare it with the following 4 state-of-the-art methods: BPR(rendle2012bpr, ), a classic pairwise collaborative filtering algorithm optimized by maximizing the difference between the probability of a user interacting with positive samples and negative samples. DiffNet(wu2019neural, ), a graph-based social recommendation method that generates final user representations by integrating user representation vectors from both the user-item interaction graph and the social graph. LightGCN(he2020lightgcn, ), a widely applied model based on NGCF that further simplifies the model structure. SocialLGN(liao2022sociallgn, ), that is currently the most advanced social recommendation graph model based on LightGCN, and GBSR (yang2024graph, ), which uses the information bottleneck principle to denoise social graphs and improve recommendation performance.